OpenStudy (anonymous):

sin(90-theta)=0.574 how do I solve this? Will give medal to best help.

OpenStudy (amistre64):

most likely with an identity. what is sine, written as cosine, look like? they are complements after all

OpenStudy (amistre64):



OpenStudy (amistre64):

90-a = -(a-90) which matches your problem better

OpenStudy (anonymous):

so, I dont know theta and am trying to solve for it, how do I do that?

OpenStudy (amistre64):

well -sin(90-theta) = -0.574 sin(theta-90) = -0.574 cos(theta) = -0.574 theta = cos^-1 (-0.574 )

OpenStudy (amistre64):

this is one approach

OpenStudy (amistre64):

ive most likely read the problem a little off in my head

OpenStudy (amistre64):

sin(90-theta)=0.574 we already have a 90-theta, i was reading it backwards .. so, sin(90-theta) = cos(theta)

OpenStudy (anonymous):

I have to find theta so that resultant force of Fa and Fb is directed horizontally to the right. I have to find magnitude of Fr

OpenStudy (anonymous):

\[\sin \left( 90-\theta \right)=\cos \theta=0.574, \theta =\cos^{-1} \left( 0.574 \right)=?\]

OpenStudy (anonymous):

I've solved everything out to this point

OpenStudy (amistre64):

 sin(90-a) = y/r cos(a) = y/r